RDJPGCOM(1) RDJPGCOM(1) NAME rdjpgcom−displaytextcommentsfromaJPEGfile SYNOPSIS rdjpgcom[−verbose][filename] DESCRIPTION rdjpgcomreadsthenamedJPEG/JFIFfile,orthestandardinputifnofileisnamed,andprints anytextcommentsfoundinthefileonthestandardoutput. TheJPEGstandardallows"comment"(COM)blockstooccurwithinaJPEGfile.Althoughthe standarddoesn’tactuallydefinewhatCOMblocksarefor,theyarewidelyusedtoholduser-sup-pliedtextstrings.Thisletsyouaddannotations,titles,indexterms,etctoyourJPEGfiles,and laterretrievethemastext.COMblocksdonotinterferewiththeimagestoredintheJPEGfile. ThemaximumsizeofaCOMblockis64K,butyoucanhaveasmanyofthemasyoulikeinone JPEGfile. OPTIONS −verbose CausesrdjpgcomtoalsodisplaytheJPEGimagedimensions. Switchnamesmaybeabbreviated,andarenotcasesensitive. HINTS rdjpgcomdoesnotdependontheIJGJPEGlibrary.ItssourcecodeisintendedasanillustrationoftheminimumamountofcoderequiredtoparseaJPEGfileheadercorrectly. In−verbosemode,rdjpgcomwillalsoattempttoprintthecontentsofany"APP12"markers astext.SomedigitalcamerasproduceAPP12markerscontainingusefultextualinformation.If youlike,youcanmodifythesourcecodetoprintotherAPPnmarkertypesaswell. SEEALSO cjpeg(1),djpeg(1),jpegtran(1),wrjpgcom(1) AUTHOR IndependentJPEGGroup 11October1997 1